OpenStudy (anonymous):

Let f(x) = 3x + 2 and g(x) =x - 3. Find f(x) - g(x). A. 2x + 5 B. 2x - 1 C. 4x - 1 D. 2x - 5

OpenStudy (igreen):

f(x) = 3x + 2 g(x) = x - 3 f(x) - g(x) 3x + 2 - x - 3 Combine like terms: 3x - x 2 - 3

OpenStudy (anonymous):

-1

OpenStudy (uri):

@iGreen Stop telling him again and again, Now next question you will do it yourself @Tazmaniadevil and we will just check.

OpenStudy (anonymous):

and -3x right

OpenStudy (uri):

3x-x=2x

OpenStudy (igreen):

Good idea. @uri

OpenStudy (anonymous):

so 2x- 1right

OpenStudy (igreen):

You got it.

OpenStudy (uri):

Right.

OpenStudy (igreen):

@Tazmaniadevil Give the medal to @uri

OpenStudy (anonymous):

thnx again lol